With a tailwind, a jet flew 2000 miles in 4 hours. The jets return trip against the same wind required 5 hours.
Find the jets speed and the wind speed.

Answer by mananth(16946)

With a tailwind, a jet flew 2000 miles in 4 hours. The jets return trip against the same wind required 5 hours.
Find the jets speed and the wind speed.
..
plane speed = xmph
current speed = y mph
..
with tailwind speed = x+y
against wind speed = x-y
..
2000/(x+y)=4
4(x+y)=2000
4x+4y=2000
/4
x+y =500 ......................1
...
2000/(x-y)=5
5(x-y)=2000
5x-5y = 2000
/5
x-y = 400......................2
add 1 & 2
2x=900
/2
x= 450 mph .... speed of plane in still air
..
x+y =500
x+450=500
x=500-450
x=50 mph ... speed of wind
...
m.ananth@hotmail.ca